# Convergence of the QuickVal Residual

### v1

@@ Abstract
$R_{n}:=\frac{S_{n}}{n}-S.$
The residual is of natural interest, especially in light of the previous analogous work on the sorting algorithm QuickSort [1, 26, 13, 16, 28]. In the case $\alpha=0$ of QuickMin with unit cost per key-comparison, we are able to calculate—à la Bindjeme and Fill for QuickSort [1]—the exact (and asymptotic) $L^{2}$-norm of the residual. We take the result as motivation for the scaling factor $\sqrt{n}$ for the QuickVal residual for general population quantiles and for general cost. We then prove in general (under mild conditions on the cost function) that $\sqrt{n}\,R_{n}$ converges in law to a scale-mixture of centered Gaussians, and we also prove convergence of moments.
@@ 1. Introduction
Parts of Sections 12 are repeated nearly verbatim, for the convenience of the reader and with permission of the publisher, from [11].
@@ 2.2. Tree of seeds and the QuickSelect tree processes
We efine the $n^{\rm th}$ QuickSelect tree process as the binary-tree-indexed stochastic process $S_{n}=(S_{n,\theta})_{\theta\in\Theta}$ and the limit QuickSelect tree process (so called in light of [11, Master Theorem 4.1]) by $I=(I_{\theta})_{\theta\in\Theta}$.
@@ 3. Exact $L^{2}$ Asymptotics for QuickMin Residual
where $\mu_{n}:=\ \mathbb{E}K_{n}=2(n-H_{n})$ for each $n$ [22]. A consequence of [19, Theorem 1] is that $K_{n}/n{\stackrel{{\scriptstyle{\mathcal{L}}}}{{\longrightarrow}}}D$, where $D{\,\stackrel{{\scriptstyle{\mathcal{L}}}}{{=}}\,}\sum_{k=0}^{\infty}\prod_{j=0}^{k}U_{j}$ has a Dickman distribution [19], with $\mu:=\mathbb{E}D=2$ (here $U_{0}:=1$). (Note that [19] refers to $D-1$ as having a Dickman distribution; we ignore this distinction.) Applying [7, Theorems 3.1 and 3.4] to the special case of QuickMin using key-comparisons costs yields the stronger result that $Y_{n}$ converges to a limit random variable $Y$ in $L^{p}$ for any $p\geq 1$ and almost surely. We can then set
@@
$I_{n}:=\#\{1<i\leq n:U_{i}<U_{1}\}$ (3.3)
(i.e., the number of keys that fall into the left subtree of the QuickSelect seed process). To begin the derivation, note that $K_{n}=n-1+\widetilde{K}_{I_{n}}$, where $\widetilde{K}_{I_{n}}$ is the key-comparisons cost for QuickMin applied to the left subtree. Note also that the same equation holds as equality in law if the process $\widetilde{K}$ has the same distribution as the process $K$ and is independent of $I_{n}$. We also have $D=1+U\widetilde{D}$ with $U:=U_{1}$ and $\widetilde{D}{\,\stackrel{{\scriptstyle{\mathcal{L}}}}{{=}}\,}D$ independent.
@@
$Y_{n}-Y$ $=\frac{n-1+K_{I_{n}}-\mu_{n}}{n+1}-\left(1+\ U\widetilde{D}-\mu\right)$
$=\left(\frac{I_{n}+1}{n+1}\right)Y_{n,0}-UY^{(0)}+\frac{n-1}{n+1}-1+\frac{\mu_{I_{n}}-\mu_{n}}{n+1}-U\mu+\mu$
$=\left(\frac{I_{n}+1}{n+1}\right)Y_{n,0}-UY^{(0)}+C_{n}(I_{n})\frac{n}{n+1}-C(U),$ (3.5)
@@
$W_{1}:=\frac{I_{n}+1}{n+1}Y_{n,0}-UY^{(0)},\quad W_{2}:=C_{n}(I_{n})\frac{n}{n+1}-C(U).$
Conditionally given $I_{n}$ and $U$, the random variable $W_{2}$ is constant and $W_{1}$ has mean zero, so
@@
$Z_{1}:=\frac{I_{n}+1}{n+1}\left(Y_{n,0}-Y^{(0)}\right),\quad Z_{2}:=\left(\frac{I_{n}+1}{n+1}-U\right)Y^{(0)},$
@@
For the cross term $\mathbb{E}(Z_{1}Z_{2})$, conditionally given $I_{n}$ the random variable $U$ is distributed Beta$(I_{n}+1,n-I_{n})$. Therefore,
$\mathbb{E}(Z_{1}Z_{2})=\mathbb{E}\left\{\mathbb{E}\left[\left.Z_{1}\left(\frac{I_{n}+1}{n+1}-U\right)Y^{(0)}\right|I_{n},Y_{n,0},Y^{(0)}\right]\right\}=0.$
@@
The conditional joint distribution of the process $(Y_{n,0})_{n\geq 0}$ and the random variable $Y^{(0)}$ given $I_{n}$ is the conditional joint distribution of the process $(Y^{*}_{I_{n}})_{n\geq 0}$ and the random variable $Y^{*}$ given $I_{n}$, where the process $(Y^{*}_{n})$ and the random variable $Y^{*}$ are independent of $I_{n}$ and have (unconditionally) the same joint distribution as the process $(Y_{n})$ and the random variable $Y$.
@@
$\mathbb{E}Z_{1}^{2}=\mathbb{E}\left[\left(\frac{I_{n}+1}{n+1}\right)^{2}\left(Y_{I_{n}}^{*}-Y^{*}\right)^{2}\right].$
Since $I_{n}\sim\text{unif}\left\{0,1,2,\ldots,n-1\right\}$, conditioning on $I_{n}$ gives
@@
Finally, consider the term $\mathbb{E}Z_{2}^{2}$. Since $Y^{(0)}$ is independent of $I_{n}$ and $U$, we have
$\mathbb{E}Z_{2}^{2}={\mathbb{E}Y^{(0)}}^{2}\,\mathbb{E}\left(\frac{I_{n}+1}{n+1}-U\right)^{2}.$
Recall that $I_{n}\sim\text{unif}\left\{0,1,\ldots,n-1\right\}$ and that conditionally given $I_{n}$, we have $U\sim\text{Beta}(I_{n}+1,n-I_{n})$; therefore,
$\mathbb{E}\left(\frac{I_{n}+1}{n+1}-U\right)^{2}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{(k+1)(n-k)}{(n+1)^{2}(n+2)}=\frac{1}{6(n+1)}.$ (3.7)
@@
We will several times invoke the following trivial result, proved by reversing the order of summation, in the sequel.
Lemma 3.4.
Let $(a_{k})_{k\geq 1}$ be a sequence of real numbers, then
$\sum_{m=0}^{n-1}\sum_{k=1}^{m}a_{k}=\sum_{k=1}^{n-1}(n-k)a_{k}.$
@@
$\mathbb{E}W_{2}^{2}=\frac{2}{3(n+1)}+\left(\frac{2}{n+1}\right)^{2}\left(1-\frac{1}{n}H_{n}\right).$
@@
$W_{2}$ $=C_{n}(I_{n})\frac{n}{n+1}-C(U)$
$=\frac{n}{n+1}\left[\frac{1}{n}\left(n-1+2(I_{n}-H_{I_{n}})-2(n-H_{n})\right)\right]+1-2U$
$=\frac{1}{n+1}\left[2(I_{n}-H_{I_{n}})-n+2H_{n}-1\right]+1-2U$
$=2\left(\frac{I_{n}+1}{n+1}-U\right)-\frac{2}{n+1}+\frac{1}{n+1}\left(-n+2H_{n}-2H_{I_{n}}-1\right)+1$
$=2\left(\frac{I_{n}+1}{n+1}-U\right)+\frac{2}{n+1}\left(H_{n}-H_{I_{n}}-1\right).$ (3.8)
@@
$\mathbb{E}W_{2}^{2}=$ $4\mathbb{E}\left(\frac{I_{n}+1}{n+1}-U\right)^{2}+\mathbb{E}\left[\left(\frac{2}{n+1}\right)^{2}\left(H_{n}-H_{I_{n}}-1\right)^{2}\right]$
$\quad+4\mathbb{E}\left[\left(\frac{I_{n}+1}{n+1}-U\right)\left(\frac{2}{n+1}\left(H_{n}-H_{I_{n}}-1\right)\right)\right].$
Recall that conditionally given $I_{n}$ we have $U\sim\text{Beta}(I_{n}+1,n-I_{n})$, which implies that the cross term vanishes; therefore, it suffices to consider the two squared terms. From (3.7) we know that
$4\mathbb{E}\left(\frac{I_{n}+1}{n+1}-U\right)^{2}=\frac{2}{3(n+1)}.$
We are now prepared to treat the final term
$L_{n}$ $:=\mathbb{E}\left[\left(\frac{2}{n+1}\right)^{2}\left(H_{n}-H_{I_{n}}-1\right)^{2}\right]$
$=\left(\frac{2}{n+1}\right)^{2}\left[\left(H_{n}-1\right)^{2}-2(H_{n}-1)\mathbb{E}H_{I_{n}}+\mathbb{E}H_{I_{n}}^{2}\right].$ (3.9)
Since $I_{n}\sim\text{unif}\{0,1,\ldots,n-1\}$, we have
$\mathbb{E}H_{I_{n}}=\frac{1}{n}\sum_{m=0}^{n-1}\sum_{k=1}^{m}\frac{1}{k}=\frac{1}{n}\sum_{k=1}^{n-1}\frac{n-k}{k}=H_{n-1}-\frac{n-1}{n},$ (3.10)
where the second equality follows from Lemma 3.4. It remains to consider
$\mathbb{E}H_{I_{n}}^{2}$ $=\sum_{m=0}^{n-1}\frac{1}{n}\left(\sum_{k=1}^{m}\frac{1}{k}\right)^{2}=\frac{1}{n}\sum_{m=0}^{n-1}\left[-\sum_{k=1}^{m}\frac{1}{k^{2}}+2\sum_{j=1}^{m}\sum_{i=1}^{j}\frac{1}{ij}\right]$
$=\frac{1}{n}\left[-\sum_{j=1}^{n-1}\frac{n-j}{j^{2}}+2\sum_{m=0}^{n-1}\sum_{j=1}^{m}\sum_{i=1}^{j}\frac{1}{ij}\right]$
$=-H_{n-1}^{(2)}+\frac{1}{n}H_{n-1}+\frac{2}{n}\sum_{m=0}^{n-1}\sum_{j=1}^{m}\sum_{i=1}^{j}\frac{1}{ij},$
where the penultimate equality follows from Lemma 3.4. Again appealing to Lemma 3.4 gives
$\frac{1}{n}\sum_{m=0}^{n-1}\sum_{j=1}^{m}\sum_{i=1}^{j}\frac{1}{ij}$ $=\frac{1}{n}\sum_{m=0}^{n-1}\sum_{j=1}^{m}\frac{1}{j}H_{j}=\frac{1}{n}\sum_{k=1}^{n-1}\frac{n-k}{k}H_{k}$
$=\sum_{k=1}^{n-1}\left(\frac{1}{k}H_{k}-\frac{1}{n}H_{k}\right)=\sum_{k=1}^{n-1}\frac{1}{k}H_{k}-\frac{1}{n}\sum_{k=1}^{n-1}\sum_{j=1}^{k}\frac{1}{j}$
$=\sum_{k=1}^{n-1}\frac{1}{k}H_{k}-\left(H_{n-1}-\frac{n-1}{n}\right).$
Plugging the last two results and (3.10) into (3.9) gives
$L_{n}$ $=\left(\frac{2}{n+1}\right)^{2}\left[\left(H_{n}-1\right)^{2}-2(H_{n}-1)\left(H_{n-1}-\frac{n-1}{n}\right)\right.$
$\qquad\qquad\qquad\left.-H_{n-1}^{(2)}+\frac{1}{n}H_{n-1}-2\left(H_{n-1}-\frac{n-1}{n}-\sum_{k=1}^{n-1}\frac{1}{k}H_{k}\right)\right].$
Note that
$H_{n-1}^{(2)}=\sum_{k=1}^{n-1}\frac{1}{k^{2}},$
so that
$-H_{n-1}^{(2)}+2\sum_{k=1}^{n-1}\sum_{j=1}^{k}\frac{1}{kj}=\left(\sum_{k=1}^{n-1}\frac{1}{k}\right)^{2}.$
This allows us to simplify the previous expression:
$L_{n}=$ $\left(\frac{2}{n+1}\right)^{2}\left[\left(H_{n}-1\right)^{2}-2(H_{n}-1)\left(H_{n-1}-\frac{n-1}{n}\right)\right.$
$\qquad\qquad\qquad\left.+H_{n-1}^{2}+\frac{1}{n}H_{n-1}-2\left(H_{n-1}-\frac{n-1}{n}\right)\right].$
After applying some algebraic manipulation, we can simplify the above expression:
$\left(\frac{n+1}{2}\right)^{2}L_{n}$ $=H_{n}^{2}-2H_{n}+1-2\left[H_{n}H_{n-1}-H_{n-1}+\left(\frac{1}{n}-1\right)H_{n}-\left(\frac{1}{n}-1\right)\right]$
$\qquad\ \ {}+H_{n-1}^{2}+\frac{1}{n}H_{n-1}-2H_{n-1}-2\left(\frac{1}{n}-1\right)$
$=\left(H_{n}^{2}-2H_{n}H_{n-1}+H_{n-1}^{2}\right)-2H_{n}+2H_{n-1}-2\left(\frac{1}{n}-1\right)H_{n}$
$\qquad+\frac{1}{n}H_{n-1}-2H_{n-1}+1+2\left(\frac{1}{n}-1\right)-2\left(\frac{1}{n}-1\right)$
$=\left(H_{n}-H_{n-1}\right)^{2}-2H_{n}+2\left(H_{n}-\frac{1}{n}\right)-2\left(\frac{1}{n}-1\right)H_{n}$
$\qquad+\frac{1}{n}\left(H_{n}-\frac{1}{n}\right)-2\left(H_{n}-\frac{1}{n}\right)+1$
$=\frac{1}{n^{2}}-\frac{2}{n}-2\left(\frac{1}{n}-1\right)H_{n}+\frac{1}{n}H_{n}-\frac{1}{n^{2}}-2H_{n}+\frac{2}{n}+1$
$=-\frac{2}{n}H_{n}+\frac{1}{n}H_{n}+2H_{n}-2H_{n}+1$
$=-\frac{1}{n}H_{n}+1.$
@@
$\mathbb{E}W_{2}^{2}=\frac{2}{3(n+1)}+\left(\frac{2}{n+1}\right)^{2}\left(1-\frac{1}{n}H_{n}\right),$

### v2

@@ Abstract
$\rho_{n}:=\frac{S_{n}}{n}-S.$
The residual is of natural interest, especially in light of the previous analogous work on the sorting algorithm QuickSort [1, 26, 13, 16, 28]. In the case $\alpha=0$ of QuickMin with unit cost per key-comparison, we are able to calculate—à la Bindjeme and Fill for QuickSort [1]—the exact (and asymptotic) $L^{2}$-norm of the residual. We take the result as motivation for the scaling factor $\sqrt{n}$ for the QuickVal residual for general population quantiles and for general cost. We then prove in general (under mild conditions on the cost function) that $\sqrt{n}\,\rho_{n}$ converges in law to a scale-mixture of centered Gaussians, and we also prove convergence of moments.
@@ 1. Introduction
Parts of Sections 12 are repeated nearly verbatim, for the convenience of the reader and with permission of the publisher, from [11]. Note, however, that we have updated the literature review, perhaps most notably including an excellent sequel to this paper, namely, [20]; see especially Remark 1.2.
@@
To motivate the reader, here is a fairly easily understood instance of our main Theorem 5.1. Suppose that keys arrive as i.i.d. uniform$(0,1)$ random variables, and suppose that cost is measured classically as the number of key comparisons. Using QuickVal to search for population quantile $\alpha\in[0,1]$, suppose for each $k\geq 0$ that the search has been narrowed to the (random) interval $(L_{k},R_{k})$ after $k$ steps of the algorithm have been carried out; in particular, $L_{0}=0$ and $R_{0}=1$. Let $I_{k}:=R_{k}-L_{k}$. Then, as shown in the proof of Theorem 5.1, the random series (of positive terms) in the expressions
$\sigma_{\infty}^{2}$ $=\sum_{k=1}^{\infty}(I_{k-1}-I_{k-1}^{2})+2\sum_{\ell=1}^{\infty}\sum_{k=1}^{\ell-1}(I_{\ell-1}-I_{k-1}I_{\ell-1})$
$=\sum_{k=1}^{\infty}I_{k}(1-I_{k})+2\sum_{\ell=2}^{\infty}I_{\ell}\sum_{k=1}^{\ell-1}(1-I_{k})$
converge with probability one, and the residual $\rho_{n}$ given by (1.1) converges in distribution to $\sigma_{\infty}Z$, where $Z$ has a standard normal distribution and is independent of $\sigma_{\infty}$.
@@ 2.2. Tree of seeds and the QuickSelect tree processes
We define the $n^{\rm th}$ QuickSelect tree process as the binary-tree-indexed stochastic process $S_{n}=(S_{n,\theta})_{\theta\in\Theta}$ and the limit QuickSelect tree process (so called in light of [11, Master Theorem 4.1]) by $I=(I_{\theta})_{\theta\in\Theta}$.
@@ 3. Exact $L^{2}$ Asymptotics for QuickMin Residual
where $\mu_{n}:=\ \mathbb{E}K_{n}=2(n-H_{n})$ for each $n$ [22]. A consequence of [19, Theorem 1] is that $K_{n}/n{\stackrel{{\scriptstyle{\mathcal{L}}}}{{\longrightarrow}}}D$, where $D{\,\stackrel{{\scriptstyle{\mathcal{L}}}}{{=}}\,}\sum_{k=0}^{\infty}\prod_{j=0}^{k}U_{j}$ has a Dickman distribution [19], with $\mu:=\mathbb{E}D=2$ (here $U_{0}:=1$). (Note that [19] refers to $D-1$ as having a Dickman distribution; we ignore this distinction.) Applying [7, Theorems 3.1 and 3.2] to the special case of QuickMin using key-comparisons costs yields the stronger result that $Y_{n}$ converges to a limit random variable $Y$ in $L^{p}$ for any $p\geq 1$ and almost surely. We can then set
@@
$N_{n}:=\#\{1<i\leq n:U_{i}<U_{1}\}$ (3.3)
(i.e., the number of keys that fall into the left subtree of the QuickSelect seed process). To begin the derivation, note that $K_{n}=n-1+\widetilde{K}_{N_{n}}$, where $\widetilde{K}_{N_{n}}$ is the key-comparisons cost for QuickMin applied to the left subtree. Note also that the same equation holds as equality in law if the process $\widetilde{K}$ has the same distribution as the process $K$ and is independent of $N_{n}$. We also have $D=1+U\widetilde{D}$ with $U:=U_{1}$ and $\widetilde{D}{\,\stackrel{{\scriptstyle{\mathcal{L}}}}{{=}}\,}D$ independent.
@@
$Y_{n}-Y$ $=\frac{n-1+K_{N_{n}}-\mu_{n}}{n+1}-\left(1+\ U\widetilde{D}-\mu\right)$
$=\left(\frac{N_{n}+1}{n+1}\right)Y_{n,0}-UY^{(0)}+\frac{n-1}{n+1}-1+\frac{\mu_{N_{n}}-\mu_{n}}{n+1}-U\mu+\mu$
$=\left(\frac{N_{n}+1}{n+1}\right)Y_{n,0}-UY^{(0)}+C_{n}(N_{n})\frac{n}{n+1}-C(U),$ (3.5)
@@
$W_{1}:=\frac{N_{n}+1}{n+1}Y_{n,0}-UY^{(0)},\quad W_{2}:=C_{n}(N_{n})\frac{n}{n+1}-C(U).$
Conditionally given $N_{n}$ and $U$, the random variable $W_{2}$ is constant and $W_{1}$ has mean zero, so
@@
$Z_{1}:=\frac{N_{n}+1}{n+1}\left(Y_{n,0}-Y^{(0)}\right),\quad Z_{2}:=\left(\frac{N_{n}+1}{n+1}-U\right)Y^{(0)},$
@@
For the cross term $\mathbb{E}(Z_{1}Z_{2})$, conditionally given $N_{n}$ the random variable $U$ is distributed Beta$(N_{n}+1,n-N_{n})$. Therefore,
$\mathbb{E}(Z_{1}Z_{2})=\mathbb{E}\left\{\mathbb{E}\left[\left.Z_{1}\left(\frac{N_{n}+1}{n+1}-U\right)Y^{(0)}\right|N_{n},Y_{n,0},Y^{(0)}\right]\right\}=0.$
@@
The conditional joint distribution of the process $(Y_{n,0})_{n\geq 0}$ and the random variable $Y^{(0)}$ given $N_{n}$ is the conditional joint distribution of the process $(Y^{*}_{N_{n}})_{n\geq 0}$ and the random variable $Y^{*}$ given $N_{n}$, where the process $(Y^{*}_{n})$ and the random variable $Y^{*}$ are independent of $N_{n}$ and have (unconditionally) the same joint distribution as the process $(Y_{n})$ and the random variable $Y$.
@@
$\mathbb{E}Z_{1}^{2}=\mathbb{E}\left[\left(\frac{N_{n}+1}{n+1}\right)^{2}\left(Y_{N_{n}}^{*}-Y^{*}\right)^{2}\right].$
Since $N_{n}\sim\text{unif}\left\{0,1,2,\ldots,n-1\right\}$, conditioning on $N_{n}$ gives
@@
Finally, consider the term $\mathbb{E}Z_{2}^{2}$. Since $Y^{(0)}$ is independent of $N_{n}$ and $U$, we have
$\mathbb{E}Z_{2}^{2}={\mathbb{E}Y^{(0)}}^{2}\,\mathbb{E}\left(\frac{N_{n}+1}{n+1}-U\right)^{2}.$
Recall that $N_{n}\sim\text{unif}\left\{0,1,\ldots,n-1\right\}$ and that conditionally given $N_{n}$, we have $U\sim\text{Beta}(N_{n}+1,n-N_{n})$; therefore,
$\mathbb{E}\left(\frac{N_{n}+1}{n+1}-U\right)^{2}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{(k+1)(n-k)}{(n+1)^{2}(n+2)}=\frac{1}{6(n+1)}.$ (3.7)
@@
$\mathbb{E}W_{2}^{2}=\frac{2}{3(n+1)}+\left(\frac{2}{n+1}\right)^{2}\left(1-\frac{H_{n}}{n}\right).$
@@
$W_{2}$ $=C_{n}(N_{n})\frac{n}{n+1}-C(U)$
$=\frac{n}{n+1}\left[\frac{1}{n}\left(n-1+2(N_{n}-H_{N_{n}})-2(n-H_{n})\right)\right]+1-2U$
$=\frac{1}{n+1}\left[2(N_{n}-H_{N_{n}})-n+2H_{n}-1\right]+1-2U$
$=2\left(\frac{N_{n}+1}{n+1}-U\right)-\frac{2}{n+1}+\frac{1}{n+1}\left(-n+2H_{n}-2H_{N_{n}}-1\right)+1$
$=2\left(\frac{N_{n}+1}{n+1}-U\right)+\frac{2}{n+1}\left(H_{n}-H_{N_{n}}-1\right).$ (3.8)
@@
$\mathbb{E}W_{2}^{2}=$ $4\mathbb{E}\left(\frac{N_{n}+1}{n+1}-U\right)^{2}+\mathbb{E}\left[\left(\frac{2}{n+1}\right)^{2}\left(H_{n}-H_{N_{n}}-1\right)^{2}\right]$
$\quad+4\mathbb{E}\left[\left(\frac{N_{n}+1}{n+1}-U\right)\left(\frac{2}{n+1}\left(H_{n}-H_{N_{n}}-1\right)\right)\right].$
Recall that conditionally given $N_{n}$ we have $U\sim\text{Beta}(N_{n}+1,n-N_{n})$, which implies that the cross term vanishes; therefore, it suffices to consider the two squared terms. From (3.7) we know that
$4\mathbb{E}\left(\frac{N_{n}+1}{n+1}-U\right)^{2}=\frac{2}{3(n+1)}.$
We now proceed to treat the final term
$L_{n}$ $:=\mathbb{E}\left[\left(\frac{2}{n+1}\right)^{2}\left(H_{n}-H_{N_{n}}-1\right)^{2}\right]$
$=\left(\frac{2}{n+1}\right)^{2}\left[\mathbb{E}(H_{n}-H_{N_{n}})^{2}-2\mathbb{E}(H_{n}-H_{N_{n}})+1\right].$ (3.9)
We can compute the first and second moments of $H_{n}-H_{N_{n}}$ as follows. Fixing $n$, for $1\leq i\leq n$ consider the events $B_{i}:=\{N_{n}<i\}$, which satisfy $\mathbb{P}(B_{i})=\frac{i}{n}$ and $B_{i}\cap B_{j}=B_{i\wedge j}$ (where $\wedge$ denotes minimum). Note that
$H_{n}-H_{N_{n}}=\sum_{i=N_{n}+1}^{n}\frac{1}{i}=\sum_{i=1}^{n}\frac{{\bf 1}(B_{i})}{i}$
Thus
$\mathbb{E}(H_{n}-H_{N_{n}})=\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})}{i}=\sum_{i=1}^{n}\frac{1}{n}=1$
and
$\mathbb{E}(H_{n}-H_{N_{n}})^{2}$ $=\mathbb{E}\left[\sum_{i=1}^{n}\frac{{\bf 1}(B_{i})}{i}\right]^{2}=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\mathbb{P}(B_{i}\cap B_{j})}{ij}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{i\wedge j}{ij}$
$=\frac{2}{n}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{j}{ij}-\frac{1}{n}\sum_{i=1}^{n}\frac{i}{i^{2}}=2-\frac{H_{n}}{n}.$
@@
$\mathbb{E}W_{2}^{2}=\frac{2}{3(n+1)}+\left(\frac{2}{n+1}\right)^{2}\left(1-\frac{H_{n}}{n}\right),$